OpenStudy (gelow):

Aerial's grandmother gave her $5,500.00 to save for her college education. She went to the bank to open a savings account. The bank told her they had two options available. Account A will pay 5.5% simple interest until the account is closed.

OpenStudy (gelow):

Account B will pay 4.5% simple interest and if the account is left open for longer than 3 years, then at the end of the third year a bonus account will be opened with $250 that also earns 4.5% for the remainder of the time the initial account is open. When the initial account is closed, the bonus account will be closed as well and the money from the two accounts will be combined.

OpenStudy (gelow):

If Aerial is planning on leaving the money in the account for 4 years and then withdrawing all funds, then account a or b is the better choice earning her ?(18.75, 30.00, 41.25, 208.75, 220.00) more.

OpenStudy (gelow):

@TheSmartOne @lizz123 @MCLover1477 @Secret-Ninja @KyanTheDoodle

OpenStudy (anonymous):

i think.....account a maybe

OpenStudy (anonymous):

im probably wrong

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

he could help

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

ort her

OpenStudy (anonymous):

or*

OpenStudy (kyanthedoodle):

Account A = 5,802.50 Account B = 6,008.75 So I believe it's B.

OpenStudy (gelow):

ok

OpenStudy (texaschic101):

account A I = PRT P = 5500 R = 5.5% = .055 T = 4 I = (5500)(.055)(4) I = 1210 total + interest = 5500 + 1210 = 6710 ------------------------ account B I = PRT P = 5500 R = 4.5% = .045 T = 3 I = PRT I = (5500)(.045)(3) I = 742.50 5500 + 742.50 = 6242.50 I = PRT P = 250 R = 4.5% = 0.45 T = 1 I = (250)(.045)(1) I = 11.25 250 + 11.25 = 261.25 6242.50 + 261.25 = 6503.75 --------------------------- Account A - Account B = 6710 - 6503.75 = 206.25 I am getting a different answer....I am getting that Account A is a better deal

OpenStudy (gelow):

i got it it was a and 41.25

OpenStudy (texaschic101):

I got A and 206.25...which is not even an answer choice

OpenStudy (texaschic101):

@Zarkon ...can you take a look at this please....where am I messing up

OpenStudy (texaschic101):

@Preetha ..can you help

OpenStudy (gelow):

i have no clue my uncle gave me the answer sorr but i did gove you a medal

OpenStudy (texaschic101):

you didn't have to give me a medal....I didn't even give you the right answer...although I cannot see my mistake.

OpenStudy (gelow):

i gave it to you because you tried

OpenStudy (texaschic101):

thank you

OpenStudy (gelow):

no problem

OpenStudy (samara8954):

I got b and 14.25

OpenStudy (gelow):

oh 41.25 was correct sorry @Samara8954

OpenStudy (zenyu):

I know this is late, but it's B and $41.25 not a